OpenStudy (anonymous):

Help to solve this one using synthetic division? Dividend: x^5-13x^4-120x+80 Divisor: x+3

OpenStudy (anonymous):

list the coefficients and put -3 on the side 1 -13 0 0 -120 80 -3 ------------------------------------------------------------------------------------------------------

OpenStudy (anonymous):

the two zeros because you have no x^3 and no x^2 now bring down the 1 1 -13 0 0 -120 80 -3 ------------------------------------------------------------------------------------------------------ 1

OpenStudy (anonymous):

1 times -3 is -3 1 -13 0 0 -120 80 -3 -3 ------------------------------------------------------------------------------------------------------ 1

OpenStudy (anonymous):

-13-3=-16 1 -13 0 0 -120 80 -3 -3 ------------------------------------------------------------------------------------------------------ 1 -16

OpenStudy (anonymous):

-3 times -16 = 48 1 -13 0 0 -120 80 -3 -3 48 ------------------------------------------------------------------------------------------------------ 1 -16

OpenStudy (anonymous):

0 + 48 = 48 1 -13 0 0 -120 80 -3 -3 48 ------------------------------------------------------------------------------------------------------ 1 -16 48

OpenStudy (anonymous):

-3 times 48 is -144 1 -13 0 0 -120 80 -3 -3 48 -144 ------------------------------------------------------------------------------------------------------ 1 -16 48

OpenStudy (anonymous):

0 -144=-144 1 -13 0 0 -120 80 -3 -3 48 -144 ------------------------------------------------------------------------------------------------------ 1 -16 48 -144

OpenStudy (anonymous):

-3 times -144 = 432 1 -13 0 0 -120 80 -3 -3 48 -144 432 ------------------------------------------------------------------------------------------------------ 1 -16 48 -144

OpenStudy (anonymous):

-120 +432 = 312 1 -13 0 0 -120 80 -3 -3 48 -144 432 ------------------------------------------------------------------------------------------------------ 1 -16 48 -144 312

OpenStudy (anonymous):

-3 times 312 = 936 1 -13 0 0 -120 80 -3 -3 48 -144 432 936 ------------------------------------------------------------------------------------------------------ 1 -16 48 -144 312

OpenStudy (anonymous):

and finally 80 + 936= 1016 1 -13 0 0 -120 80 -3 -3 48 -144 432 936 ------------------------------------------------------------------------------------------------------ 1 -16 48 -144 312 1016

OpenStudy (anonymous):

aah made a mistake on the very last line. should be -3 times 312= -936 and 80 - 936 = -856 rest is right though

OpenStudy (anonymous):

answer is \[x^4-16x^3+48x^2-144x+312 -\frac{856}{x+3}\]